OpenStudy (anonymous):

Given the point (3, 4) and the slope of 6, find y when x = 27.

OpenStudy (anonymous):

kadkin is typing

OpenStudy (anonymous):

So it would look like 2x-3y=6

OpenStudy (anonymous):

^how did u get that equation

OpenStudy (anonymous):

so 2(27)-3y=6

OpenStudy (anonymous):

I got the equation because from google actually

OpenStudy (anonymous):

when i typed in the question online then i was reading through and it gave me this equation hold on and ill give you the webpage. http://www.algebra.com/algebra/homework/Linear-equations.faq.question.86726.html

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it gives you directions on how to do it

OpenStudy (anonymous):

himluv- if you need any help understanding how to get that equation, please let me know

OpenStudy (anonymous):

thanks for being there to explain it to him but i know you get the 3 from your x value and the 6 for your slope and the x you replace with the 27 but not the rest so you can explain :)

OpenStudy (anonymous):

^yes, you are correct. We can use y=mx+c equation. We know m=6 and the equation becomes y=6x + c. Then using the given coordinates to x and y values, we can find c. and substituting both m and c to our original equation will give the final equation you already got

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

do you want me to give the answer choices?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

are you gona type them or what hun

OpenStudy (anonymous):

122 14 141 148

OpenStudy (anonymous):

I'm afraid kadkins's equation is wrong. But the correct answer is 148. and the correct equation is y=6x-14

OpenStudy (anonymous):

my badd broskieess

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

internet can be stupidd!! :)